A boat goes 280 miles downstream in the same time it can go 210 miles upstream. The speen of the current is 9 miles per hour. Find the speed of th boat in still water.
Answer by Paul(988)

Let the speed of the boat be x
280(x-9)=210(x+9)
280x-2520=210x+1890
70x=4410
x=63

Hence, the speed of the boat is 63mph.
PAUL.